A jewelry maker made two purchases. The first purchase, which cost $220, included 50 grams of a silver alloy and 40 grams of a copper alloy. The second purchase, at the same prices, cost $140 and included 25 grams of the silver alloy and 30 grams of the copper alloy. Find the cost per gram of the silver alloy and the copper alloy.
Answer by scott8148(5880)

let s=silver, c=copper ... 50s+40c=220 ... 25s+30c=140 ... multiplying second equation by -2 and adding to first equation gives

50s+40c-2(25s+30c)=220-2(140) ... 50s+40c-50s-60c=220-280 ... -20c=-60 ... c=3

50s+40(3)=220 ... 50s+120=220 ... 50s=100 ... s=2